Recent Log Entries

HYST AREA 6-2-4 Cycles Magnets

Log Number: 4152402Logbook(s): Body: 
HYST AREA Version 6-2-4 has FAILED to cycle:

MBC2T01H -- W202B16 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T03V -- W202B16 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T02H -- W202B16 -- wa2mag -- Failed to set after 2 cycle attempts


HYST AREA Version 6-2-4 has successfully cycled the following:

Area.Transport.2T

The following magnets were REMOVED from the CYCLING LIST by the user:
MBC2T01V
MBC2T02H
MBC2T03H
MBC2T03V
MBC2T04H
MBC2T04V
MBC2T05V
MBC2T06H

HYST AREA dvl Cycles Magnets

Log Number: 4152397Logbook(s): Body: 
HYST AREA Version dvl has FAILED to cycle:

MBC2T08H -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T09H -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts

MBC2T07H -- W202B14 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T06H -- W202B14 -- wa2mag -- Failed to set after 2 cycle attempts

MBC2T01H -- W202B16 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T03H -- W202B16 -- wa2mag -- Failed to set after 2 cycle attempts

HYST AREA dvl Cycles Magnets

Log Number: 4152395Logbook(s): Body: 
HYST AREA Version dvl has FAILED to cycle:

MBC2T08H -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T08V -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T09H -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T09V -- W202B13 -- wa2mag -- Failed to set after 2 cycle attempts

MBC2T07H -- W202B14 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T07V -- W202B14 -- wa2mag -- Failed to set after 2 cycle attempts
MBC2T06H -- W202B14 -- wa2mag -- Failed to set after 2 cycle attempts

COMPLETED Update JELI to version 5-0

Log Number: 4152370Logbook(s): Body: 

I went ahead and made a test entry from every OPS and CHL machine
- one issue found - wish link had to be recreated (tcl/bin/8.6/wish -> wish8.6) for RHEL6 machines

SL tunnel flood cleanup completed

Log Number: 4152352Logbook(s): Tags: Entry Makers: friersonBody: 

Done! It is safe to walk though again. I want to give a huge thanks to INSTALL and FMM for all the help to make this happen.

Sorry for the side-pics….just tilt your head.

where to cap off chopper #2

Log Number: 4152318Logbook(s): Entry Makers: gramesBody: 

Fig. 1 shows the beamline downstream of chopper #2.

Checking with Matt, we propose to remove the BPM+bellows, then blank off upstream nipple (to pumpdown) and downstream cross (to keep clean and protect). The corrector and BPM cables will be removed in advance. The earth coil might still be there.

Gun2 photocathodes (hardcopy logbook)

Log Number: 4152313Logbook(s): Entry Makers: gramesBody: 

Looking at the Gun2 logbook (fig. 1) I confirm the linked entry, i.e. what we have in Gun2 today...

HV cathode => puck #9 - SLSP-5756-4 (CREX/PREX-II)

SM#1 => puck #32 bulk used for HV conditioning
SM#2 => puck #3 (furthest heater) - SLSP-5247-4 (QWeak)
SM#2 => puck #36 - SLSP-DBR-75102-2 (unused)

#9 (nearest heater) - SLSP-5756-4 (CREX/PREX-II)

HYST AREA dvl Cycles Magnets

Log Number: 4152307Logbook(s): Body: 
Unable to connect to the following IOCS:
sb4mag

HYST AREA Version dvl has successfully cycled the following:

BS04B10

The following magnets were REMOVED from the CYCLING LIST by the user:
MQD1L04
MCA1L01
MBT1L00H
MBT1L00V
MBT1L01H
MCB1L01
MCB1L01A
MQB1L08
MQB1L09
MQB1L10
MQB1L11
MQB1L12
MQB1L13
MQB1L14
MQB1L15
MQB1L16
MQB1L17
MQB1L18
MQB1L19
MQB1L20
MQB1L21
MQB1L22
MQB1L23
MQB1L24
MQB1L25
Total Time: 0.033 mins

HYST AREA dvl Cycles Magnets

Log Number: 4152306Logbook(s): Body: 
Unable to connect to the following IOCS:
sb4mag

HYST AREA Version dvl has successfully cycled the following:

BS04B10

The following magnets were REMOVED from the CYCLING LIST by the user:
MQD1L04
MCA1L01
MBT1L00H
MBT1L00V
MBT1L01H
MCB1L01
MCB1L01A
MQB1L08
MQB1L09
MQB1L10
MQB1L11
MQB1L12
MQB1L13
MQB1L14
MQB1L15
MQB1L16
MQB1L17
MQB1L18
MQB1L19
MQB1L20
MQB1L21
MQB1L22
MQB1L23
MQB1L24
MQB1L25
Total Time: 0.033 mins

SAD Service Building status

Log Number: 4152302Logbook(s): Tags: Body: 

RCG-1 lock applied to in-line dump. All service building RCA postings have been removed. East and West ARC service building RA postings have been flipped to the de-posted state and all HRA roof signage removed. North and South LINAC C100 zone klystron RA signage flipped to access allowed state. Fences around north and south He vents and E1 fence de-posted. All RCG-1 locks removed from ARC service building racks and E1 fence.

Box PS LOTO

Log Number: 4152299Logbook(s): Body: 

Kevin and I performed LOTO on all South Access power supplies (ARC1, ARC3, ARC5, ARC7, ARC9 and RSEP9), Even Dogleg supplies (DOG2, DOG4, DOG6 and DOG8), XSEPT Supplies (XSEP2,T XSEP4T, XSEP6T, XSEP8T, XSEPAT). YA's were also powered down and turned off.

COMMENT ON Injector Beamline Repairs: Winter 2022 SAD

Log Number: 4152296Logbook(s): Body: 

Update March 2023 - anticipated dates for each step

Monday & Tuesday, March 20-21: EES de-energize & lock out beamline magnets, chopper RF power

Wednesday, March 23: Marcy & Phil power down all affected ion pump power supplies including UHV supply for 0I02.
Vent beamline from VBV0I02 through VBV0I07

Replace harp which is long enough to interfer with top hat: harp checked in lab before install (Tony's help for harp motion check)
Return top hat to viewer IGY0I02 so it doesn't charge up and steer beam
Check out to ensure no interference.

Operator SIGN IN

Log Number: 4152290Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Thomas
DATE:             03/20/2023
SHIFT:               Day

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Pages